OpenStudy (anonymous):

is the following statement true? 3(a/b) = (3a)/(3b)

OpenStudy (zepp):

This statement is absolutely false.

OpenStudy (anonymous):

no its FALSE

OpenStudy (anonymous):

3 cancels out with each other and you are left with a/b

OpenStudy (zepp):

\(\large 3(\frac{a}{b})=\frac{3a}{b}\)

OpenStudy (anonymous):

\[3\frac{a}{b} = \frac{3 \times a}{b}\] So, it is false...

OpenStudy (anonymous):

thanks everyone:)..ya"ll are life savors<3:)